# SenseCF: LLM-Prompted Counterfactuals for Intervention and Sensor Data Augmentation

### v1

@@ I Introduction and Related Work
Recent work highlights LLMs’ innate counterfactual reasoning capabilities without fine-tuning [2, 3], yet their use in structured, multimodal health data remains underexplored. Importantly, CFs not only enhance interpretability but can also serve as data augmenters, introducing label-flipping samples to strengthen models—particularly in imbalanced medical datasets. For data augmentation, CFs enhance robustness by introducing label-flipping variations while preserving data distributions. Optimization methods show promise in medical/low-data contexts but struggle with categorical coherence—a gap addressed by LLMs’ semantic understanding [4].
@@ II Methods
We represent our input data as a set of tuples $(x_{i},y_{i})$ , where $x_{i}\in X$ is a feature vector representing either clinical or physiological data and $y_{i}\in\{0,1\}$ denotes the ground truth label. The trained predictive model $f(\cdot)$ outputs predictions $\hat{y_{i}}=f(x_{i})$. Our preprocessing stage transforms raw data into structured, tabular feature vectors suitable for prompting LLM.
@@ II-A Counterfactual Generation
$x^{\prime}_{i}=\text{LLM}(x_{i},\text{prompt}),\quad\text{subject to }f(x^{\prime}_{i})\neq f(x_{i})\vspace{-1.2mm}$
@@ II-B Intervention and Data Augmentation
$X_{\text{aug}}=X\cup\{(x^{\prime}_{i},y^{\prime}_{i})\mid f(x^{\prime}_{i})\neq f(x_{i})\}\vspace{-1mm}$
@@ III-A Data
1. The AI‑READI data, part of an NIH-funded initiative, combines rich data sources—like vitals, ECG signals, continuous glucose monitoring (CGM), physical activity, and eye imaging—collected from 1,067 participants, both diabetic and non-diabetic. In this analysis, we used twelve features to design counterfactual interventions for transitioning individuals from high stress ($y=1$) to moderate stress levels ($y=0$). Of the twelve features, we considered age, sex, and medication as immutable features; the rest are either raw or derived features from the sleep and daily glucose information collected using sensors (Garmin Vivosmart 5 and Dexcom G6).
2. Modified Heart Disease dataset from [5] has $918$ instances. We used four categorical and five continuous features to map whether an individual will have a heart disease $(y=1)$ or not $(y=0)$ and make interventions based on five features. A subset of the features, including fasting blood sugar level, resting blood pressure, resting ECG, and exercise-induced S-T depression, are derived using wearable sensors.
@@ III-B Baselines
DiCE [6] identifies a set of CFs by optimizing for proximity, diversity and sparsity. CFNOW [7] searches an optimal point close to the factual point where the classification differs from the original. CFNOW performs greedy optimization for metrics like speed, coverage, distance, and sparsity. NICE [4] CFs are not necessarily adversarial data points but nearby instances in the data that reflects the desired outcome.
@@ III-C Evaluation Metrics
$\vspace{-1.5mm}\textit{validity}=\frac{\#|f(X_{T}^{*})\neq f(X_{T})|}{\|CF\|}\vspace{-1.5mm}$
@@
$\vspace{-3mm}\textit{sparsity}=\frac{\sum_{X_{T}^{*}\in CF}\sum_{i=1}^{d}\mathbbm{1}(x_{T}^{*i}\neq x_{T}^{i})}{\|CF\|}\vspace{-1mm}$ (1)
@@
$\vspace{-1mm}\textit{plausibility}=\frac{\sum_{X_{T}^{*}\in CF}\mathbbm{1}(\text{dist}(X_{T}^{*})\subseteq\text{dist}(X))}{\|CF\|}\vspace{-1.5mm}$
@@ IV-A Intervention
A 81-year-old patient labeled as “stressed” showed low deep sleep (30.1%), moderate REM (15.4%), high blood glucose (210.8 mg/dL), and limited activity (5.95 steps). Stress level was high (85.25), with poor glucose control (TIR: 12.5%, 1 hyper event).   The LLM suggests increasing deep sleep to $\uparrow$35and REM sleep to $\uparrow$20%, which could help reduce physiological and emotional stress. It also recommends lowering blood glucose from 210.8 to $\downarrow$180 mg/dL, aligning with better metabolic control. These changes reflect clinically actionable strategies such as sleep hygiene improvement and tighter glucose management.
@@
Figure 4 shows the diversity of the CFs from different methods across all the mutable features. SenseCF-generated CFs exhibit least diversity for all features except that the 3-shot variant shows higher diversity in Average Step counts.

### v2

@@ I Introduction and Related Work
Recent work highlights LLMs’ innate counterfactual reasoning capabilities without fine-tuning [2, 3], yet their use in structured, multimodal health data remains underexplored. Importantly, CFs not only enhance interpretability but can also serve as data augmenters, introducing label-flipping samples to strengthen models—particularly in imbalanced medical datasets. For data augmentation, CFs enhance robustness by introducing label-flipping variations while preserving data distributions. Optimization methods show promise in medical or low-data contexts but struggle with categorical coherence—a gap addressed by LLMs’ semantic understanding [4].
@@ II Methods
We represent our input data as a set of tuples $(x_{i},y_{i})$ , where $x_{i}\in X$ is a feature vector representing either clinical or physiological data and $y_{i}\in\{0,1\}$ denotes the ground truth label. The trained predictive model $f(\cdot)$ outputs predictions $\hat{y_{i}}=f(x_{i})$. Our preprocessing stage transforms raw data into structured, tabular feature vectors suitable for prompting the LLM.
@@ II-A Counterfactual Generation
$x^{\prime}_{i}=\text{LLM}(x_{i},\text{prompt}),\quad\text{subject to }f(x^{\prime}_{i})\neq f(x_{i})\vskip-3.41432pt$
@@ II-B Intervention and Data Augmentation
$X_{\text{aug}}=X\cup\{(x^{\prime}_{i},y^{\prime}_{i})\mid f(x^{\prime}_{i})\neq f(x_{i})\}\vskip-2.84526pt$
@@ III-A Data
1. The AI‑READI data, part of an NIH-funded initiative, combines rich data sources—like vitals, ECG signals, continuous glucose monitoring (CGM), physical activity, and eye imaging—collected from 1,067 participants, including both diabetic and non-diabetic. In this analysis, we used twelve features to design counterfactual interventions for transitioning individuals from high stress ($y=1$) to moderate stress levels ($y=0$). Of the twelve features, we considered age, sex, and medication as immutable features; the rest are either raw or derived features from the sleep and daily glucose information collected using sensors (Garmin Vivosmart 5 and Dexcom G6).
2. Modified Heart Disease dataset from [5] has $918$ instances. We used four categorical and five continuous features to map whether an individual will have a heart disease $(y=1)$ or not $(y=0)$ and design interventions based on five features. A subset of the features, including fasting blood sugar level, resting blood pressure, resting ECG, and exercise-induced S-T depression, are derived using wearable sensors.
@@ III-B Baselines
DiCE [6] identifies a set of CFs by optimizing for proximity, diversity and sparsity. CFNOW [7] searches an optimal point close to the factual point where the classification differs from the original. CFNOW performs greedy optimization for metrics like speed, coverage, distance, and sparsity. NICE [4] CFs are not necessarily adversarial data points but nearby instances in the data that reflect the desired outcome.
@@ III-C Evaluation Metrics
$\vskip-4.2679pt\textit{validity}=\frac{\#|f(X_{T}^{*})\neq f(X_{T})|}{\|CF\|}\vskip-4.2679pt$
@@
$\vskip-8.53581pt\textit{sparsity}=\frac{\sum_{X_{T}^{*}\in CF}\sum_{i=1}^{d}\mathbbm{1}(x_{T}^{*i}\neq x_{T}^{i})}{\|CF\|}\vskip-2.84526pt$ (1)
@@
$\vskip-2.84526pt\textit{plausibility}=\frac{\sum_{X_{T}^{*}\in CF}\mathbbm{1}(\text{dist}(X_{T}^{*})\subseteq\text{dist}(X))}{\|CF\|}\vskip-4.2679pt$
@@ IV-A Intervention
A 81-year-old patient labeled as “stressed” showed low deep sleep (30.1%), moderate REM (15.4%), high blood glucose (210.8 mg/dL), and limited activity (5.95 steps). Stress level was high (85.25), with poor glucose control (TIR: 12.5%, 1 hyper event).   The LLM suggests increasing deep sleep to $\uparrow$35% and REM sleep to $\uparrow$20%, which could help reduce physiological and emotional stress. It also recommends lowering blood glucose from 210.8 to $\downarrow$180 mg/dL, aligning with better metabolic control. These changes reflect clinically actionable strategies such as sleep hygiene improvement and tighter glucose management.
@@
Figure 4 shows the diversity of the CFs from different methods across all the mutable features. SenseCF-generated CFs exhibit the least diversity for all features, except that the 3-shot variant shows higher diversity in Average Step counts.